OpenStudy (anonymous):

Some people answered this question but they said it was wrong. What is the simplest fraction whose value is equal to the number / depicted on this number line?

OpenStudy (anonymous):



OpenStudy (anonymous):

could you clarify your number line a bit please? is it supposed to be 2 and 2/7 or what? Is that a 1 or an arrow...?

OpenStudy (anonymous):

it's supposed to be a slash

OpenStudy (anonymous):

okay, just fully read your question. makes sense. Is the number line supposed to be divided into 7 equal sections starting at 2 and ending at 3?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

alrighty! So since you are past 2 on the number line but before 3, you know you have 2 plus some fraction of a whole number. One unit on this number line is divided evenly in 7 parts by the 7 tick marks. That means that for each tick mark on the number line, you add 1/7 to your number Your slash / is 2 tick marks to the right from 2. That means you have 2 + 2/7 = 2 and 2/7 you can simplify this by putting both the whole number in the fraction in the same form with the same denominator the least common denominator is 7. 2 = 14/7 therefore you have 2 + 2/7 = 14/7 + 2/7 = 16/7